Hello, I have a question on this problem ::::
Write a slope-intercept equation of the line whose graph is described.
Perpendicular to the graph of 8x+y=2 ; y-intercept (0,6)
A)y=8x+6
B)y=1/8x+6
C)y=-8x+6
D)y=-1/8x+6

Answer by mananth(16946)

8 x + 1 y = 2
Find the slope of this line

1 y = -8 x + 2
Divide by 1
y = -8 x + 2
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = -8

The slope of a line perpendicular to the above line will be the negative reciprocal 1/8
Because m1*m2 =-1
The slope of the required line will be 1/8

m= 1/8 ,point ( 0 , 6 )
Find b by plugging the values of m & the point in
y=mx+b
6 = 0 + b
b= 6
m= 1/8
The required equation is y = 1/ 8 x + 6
m.ananth@hotmail.ca